Two planes are being evaluated, with the projected life of each being 10 years. The cash flow for each plane are below. Use MARR of 15%. Based on the rate of return which is the most better.

A B
First Cost $350 000 $415 000
M & O Costs 15 000 10 000
Annual Benefit 90 000 125 000
Salvage Value 35 000 45 000
Answer #1

Plane A

The present value of cash flow can be written as follows

large P = rac{A}{1+IRR}+...rac{A}{(1+IRR)^n}+rac{S}{(1+IRR)^n}

For plan A

A = 90000 - 15000 = $ 75000 per year

Salvage cost, S = $ 35000

We can calculate IRR using Trial and error method

Let us assume IRR = 20%

PA  = 75,000(PVIFA,20%,10) + 35,000(PVIF,20%,10)

= 75,000*4.192 +35,000*0.1615

= 314,400 + 5652.70

= $ 320,052.70

Now assume IRR = 10%

PA = 75,000*6.1445 + 35,000*0.3855

= 460,842.5 + 13,494.01

= $ 474,336.51

Thus the IRR of plan A lies in between 10 to 20%. We can determine it using linear interpolation technique as follows

IRRA 10(20 - 10)* 474,336.51 - 350,000 474,336.51 - 320,052.7

124,336.51 IRRA 10+(20-0) 52838

「RR 10 + 10 * 0.8059

large IRR_A = 10 + 8.059

IR RA = 18.059%

Similarly calculate for plan B

Let us assume when IRR is 20%

PB = 115,000(PVIFA,20%,10) +45,000(PVIF,20%,10)

= 115,000*4.192 + 45,000*0.1615

= $ 489,347.50

Since the initial investment is less than the calculated value thus increase the rate

Let us assume IRR = 30%

PB = 115,000*3.10 + 45,000*0.072538

= 355,527 + 3264.22

= $ 358,791.20

Linear interpolation technique as follows

489, 347.5-415,000 KB20+ (30489,347.5 358,791.2 258 70

4,347.5 IRRB- 20 + 10* 130,556.2

IRR в 20 + 5.69

IRRB = 25.69%

Both the plans can be selected since marr is less PV.

Select plan B

Please contact if having any query thank you.